OpenStudy (anonymous):

The portion of the city's trash that is recyclable is 2/3.If 3/8 of the recyclable trash is cans,what fraction of the city's trash is recyclable cans?Reduce the answer to the lowest terms. A. 24/25 of the trash B. 9/16 of the recyclable trash C. 1/4 of the recyclable trash D. 7/24 of the recyclable trash

OpenStudy (anonymous):

Help Medal rewarded!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

try multiplying the denominator and the numerator by a number until it has the same multiple as the other denominator

OpenStudy (anonymous):

like this\[\frac{ 3 }{ 8 }*\frac{ 3 }{ 3 }=\frac{ 9 }{ 24 }\] the denominators of the 2 fractions 8 and 3 are multiples of 24

OpenStudy (anonymous):

so multiply 2/3 by some multiple of 28.\[\frac{ 2 }{3 }*\frac{ x }{ x }=\frac{ x }{ 24 }\] remember, the number you're multiplying by the denominator has to be the same number to multiply the numerator

OpenStudy (er.mohd.amir):

(2/3) * (3/8) =2/8=1/4